OpenStudy (anonymous):

79 1/6, change to improper fraction.

OpenStudy (anonymous):

475/6

OpenStudy (diyadiya):

79*6+1/6=475/6

OpenStudy (anonymous):

to do this first you need to multiply 79*6=474 add the 1 474+1= 475 so now the answer is 475/6